OpenStudy (anonymous):

One 10.0 Ω resistor is wired in series with two 5.0 Ω resistors in parallel. A 9.0 V battery supplies power to the circuit. What is the total current through the circuit? A. 0.45 A B. 0.72 A C. 0.87 A D. 1.1 A

OpenStudy (michele_laino):

here is our circuit: 

OpenStudy (anonymous):

ok! what do we do? :/

OpenStudy (michele_laino):

here the total resistance or equivalent resistance is: \[{R_{TOTAL}} = 10 + \frac{5}{2} = ...Ohms\]

OpenStudy (anonymous):

ok! we get 12.5?

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

so , applying the Ohm's law, we get: \[\Large I = \frac{V}{{{R_{TOTAL}}}} = \frac{9}{{12.5}} = ...Amperes\]

OpenStudy (anonymous):

0.72! so choice B is our solution?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

yay! thanks!!:)

OpenStudy (michele_laino):

:):)